Three research departments have 10, 6, and 8 members, respectively. Each department is to select a delegate and an alternate to represent the department at a conference. In how many ways can this be done?
Answer #1

1st department (10choices for the delegate and 9 for the alternate, which has 10*9=90 )

2nd department ( 6 choices for the delegate and 5for the alternate, which has 6*5=30)

3rd department ( 8 choices for the delegate and 7for the alternate, which has 8*7= 56 )

therfore, there are 90*30*56 =151200 ways